# Integrating Reinforcement Learning and Model Predictive Control with Applications to Microgrids

### v1

@@ Abstract
This work proposes an approach that integrates reinforcement learning and model predictive control (MPC) to efficiently solve finite-horizon optimal control problems in mixed-logical dynamical systems. Optimization-based control of such systems with discrete and continuous decision variables entails the online solution of mixed-integer quadratic or linear programs, which suffer from the curse of dimensionality. Our approach aims at mitigating this issue by effectively decoupling the decision on the discrete variables and the decision on the continuous variables. Moreover, to mitigate the combinatorial growth in the number of possible actions due to the prediction horizon, we conceive the definition of decoupled Q-functions to make the learning problem more tractable. The use of reinforcement learning reduces the online optimization problem of the MPC controller from a mixed-integer linear (quadratic) program to a linear (quadratic) program, greatly reducing the computational time. Simulation experiments for a microgrid, based on real-world data, demonstrate that the proposed method significantly reduces the online computation time of the MPC approach and that it generates policies with small optimality gaps and high feasibility rates.
keywords:
reinforcement learning , model predictive control , hybrid systems , mixed-integer programming , power systems , mixed-logical dynamical systems , machine learning
[…]
[1]organization=Delft University of Technology, addressline=Melkweg 2, postcode=2628 CD, city=Delft, country=The Netherlands
@@ 1.1 Motivation
Complex infrastructure systems, such as energy, water, and transportation networks are pervasive in our modern world. Analysis and control design of such systems is very challenging due to their size and intricate behavior. These networks inherently have a combination of discrete and continuous dynamics and/or decision variables. Several modeling approaches [1, 2, 3] have considered hybrid systems to represent these important infrastructure networks. For instance, increasing energy demand and expanding decentralized energy generation by renewable sources have motivated the conception of microgrids, which are local electrical grids that can be connected to a main grid. The operation of such microgrids consists of scheduling both the power flow directions and the amount of power being exchanged between each of the entities of the microgrid. As a result, discrete and continuous decision variables have to be planned considering market conditions to reduce the operation cost.
@@
For the considered problem, model predictive control (MPC) arises as a promising technique due to its ability to handle constrained complex systems with discrete and continuous decision variables [4]. In MPC, the system model is used for optimization of the control actions over a finite prediction horizon. At each time step, the model-based optimization problem is solved, the first element of the control sequence is applied to the system, and the prediction horizon is shifted one time step ahead. Furthermore, MPC has solid theoretical foundations with regard to stability, performance, and safety [5]. However, these advantages often come at the price of intense online computational requirements, limiting the use of MPC due to hardware constraints or execution time limitations. This bottleneck is especially more pronounced in MPC for hybrid systems, where the optimizer has also to consider a sequence of discrete decision variables over the given prediction horizon. In this case, the optimization problem then becomes of mixed-integer nature, which is NP-hard [6].
@@
The most widely used technique to solve mixed-integer problems in modern solvers is branch-and-bound [7]. In the worst case, the solver has to find solutions for the relaxed problems for each possible combination of the integer decision variables. This is not scalable with respect to the number of integer variables because of the combinatorial nature of the problem. Branch-and-bound algorithms mitigate this issue by efficiently pruning branches of the search tree by estimating lower and upper bounds of the objective function and by proposing cutting planes, which reduce the feasible set of the mixed-integer programs by introducing linear inequalities as additional constraints [8]. Moreover, many expert-designed heuristics are employed for improving the search such as node selection and branch variable selection techniques to reduce computation time. One alternative to branch-and-bound for the control of hybrid systems is to pre-compute the MPC control law offline and simply evaluate this function online. For hybrid systems, the off-line computation of the explicit MPC control law via multi-parametric programming and dynamic programming was explored in [6]. However, this approach, also referred to as explicit MPC, can be only successfully applied to low-dimensional linear systems. Despite improvements in branch-and-bound and explicit MPC, solving mixed-integer problems with a significant number of integer decision variables is fundamentally difficult.
@@
Recently, supervised learning has been explored in several MPC approaches for hybrid systems to reduce the online computational time of the resulting mixed-integer programs [9, 10, 11, 12]. In essence, the aforementioned supervised learning methods have the same structure and learning setting and the main difference lies in the choice of the classifier. These works employ supervised learning to approximate the mapping from the system state to the discrete optimal solution. By applying this classifier to predict the discrete optimization variables, the mixed-integer program is then simplified to an optimization problem consisting only of real-valued variables, significantly reducing the online computational burden of the MPC controller. To build the training dataset, these methods rely on branch-and-bound to solve the control problem – a mixed-integer program – to optimality several times. As a result, the main issue is sidelined to the offline phase of the algorithm, where typically more computing resources are available. Moreover, in supervised learning the goal is to reduce the classification error, e.g. the distance between the predicted discrete sequence and the optimal discrete sequence. Even though the classification error is typically a suitable proxy for control performance, this might not necessarily be true depending on the mapping from the input sequence to the control objective function. For a more complete overview of the literature on the intersection of learning and control for hybrid systems, the reader is referred to Section 2.
@@ 1.2 Contributions
Our work aims to propose an integrated reinforcement learning and MPC method that solves mixed-integer linear and quadratic programs with low computational footprint, low optimality gap, and high feasibility rate. We build on the existing idea of decoupling the decision on the discrete and continuous variables with learning and MPC [9, 10, 11, 12]. However, we explore a novel paradigm by employing reinforcement learning in place of supervised learning to directly optimize for control performance – instead of minimizing the classification error – and to avoid the use of branch-and-bound in both the offline and online phases of the algorithm.
@@
  • 1.
    We propose a novel integrated reinforcement learning and MPC framework for control of mixed-logical dynamical systems.
  • 2.
    The Q-function is partitioned across the prediction horizon and the definition of decoupled Q-functions is conceived to make the learning problem more tractable.
  • 3.
    Simulation experiments in a microgrid system show the efficacy of the proposed approach in reducing the computational load of the MPC controller. Moreover, the comparison between the proposed approach and a method based on supervised learning reveals a trade-off: while the former outperforms in terms of feasibility, the latter outperforms in terms of optimality.
In our problem, long-term decision-making involves determining suitable sequences of discrete decision variables over a given prediction horizon, which is a combinatorial problem. To mitigate the combinatorial complexity of the optimization problem, we propose the decoupling of the Q-function across time, i.e., each time step of the prediction horizon has its own decoupled Q-function. With the proposed decoupling of the Q-function, our method effectively simplifies the learning problem as the discrete actions are decided sequentially rather than simultaneously. A similar idea was explored in [13]; however, the separation therein was not in time, but for independent actions. In our setting, future actions may depend on past actions; therefore, we use a recurrent neural network to retain past information in its internal states and recursively compute the decoupled Q-functions. A case study is used to empirically show that the reinforcement learning agent can learn performing policies with the unorthodox definition of decoupled Q-functions.
@@ 1.3 Outline
This paper is organized as follows. In Section 2, we give an overview of the literature on the intersection of learning and control for hybrid systems. Section 3 formalizes the control problem. Our novel method that integrates reinforcement learning and MPC for the control of mixed-logical dynamical systems is described in Section 4. Section 5 presents the simulation setup, the results, and the discussion. This paper ends in Section 6 with conclusions and suggestions for future work.
@@ 2 Related work
In [17] the authors approximate the state-action value function (Q-function) of time-invariant mixed-logical dynamics systems. The core of their approach is to parametrize the Q-function with Benders cuts and estimate the Q-function from a lower bound. However, the scope of application of this approach is limited because it only considers auxiliary binary variables from the mixed-logical dynamical modeling framework and it is not compatible with binary states and inputs. In [18], the authors are successful in reducing the computational burden of MPC by learning a state-dependent horizon and a state-dependent recomputation policy, i.e. the policy that decides whether the MPC optimal control problem should be recomputed at a given time step. In [19], an MPC controller is used as a function approximator of the state-action value. The authors show how the internal parameters of the controller can be tuned by policy gradient methods. An extension of the same approach to mixed-integer problems was made in [20]; however, the goal is to target performance rather than to reduce the computational cost of the optimization problem.
There is also a body of work that explores the use of reinforcement learning to jointly learn discrete and continuous policies [21, 22, 23]. In the artificial intelligence community, this problem is explored under the framework of Markov decision processes with parameterized actions. The problem is very similar to that of control of hybrid systems because the action space has discrete and continuous elements. Even though these works can provide useful insight into the parametrization and training of policies for hybrid systems, they lack the optimality and constraint satisfaction that MPC can provide.
@@ 3 Control problem
As a control algorithm, we consider MPC for its capacity to handle multivariable constrained hybrid systems. In the context of MPC, mixed-logical dynamical (MLD) systems [26] are typically used to model hybrid systems since they can be readily used to formulate open-loop finite-horizon optimal control problems [4]. The equivalence of MLD systems and other hybrid system modeling frameworks was established in [27], showing their broad applicability. When MLD systems are controlled with MPC and the cost function and the constraints are linear (quadratic), the resulting optimization problem is a mixed-integer linear (quadratic) program (MIL(Q)P). In this section, a general description of the MPC optimization problem for an MLD system is first addressed to set the stage for the formulation of the control problem as an MIL(Q)P.
@@
where $x\in\mathbb{R}^{n_{\mathrm{c}}}\times\{0,1\}^{n_{\mathrm{d}}}$ is a vector containing the continuous and discrete system states, $u\in\mathbb{R}^{m_{\mathrm{c}}}\times\{0,1\}^{m_{\mathrm{d}}}$ are the continuous and discrete inputs, $\delta\in\{0,1\}^{r_{\mathrm{d}}}$ is a vector with the auxiliary discrete variables, $z\in\mathbb{R}^{r_{\mathrm{c}}}$ is a vector with the continuous auxiliary variables arising from the MLD modeling, and $A,\ \{B_{i}\}_{i=1,2,3,5},\ \{E_{i}\}_{i=1,2,3,4,5}$ are matrices of appropriate dimensions. Note that the linear constraints may represent logical constraints, a byproduct of MLD modeling, or system operating constraints.
@@
$\begin{split}\min_{\mathbf{x}(k),\epsilon_{\mathrm{c}}(k),\epsilon_{\mathrm{d}}(k)}\ &J(\mathbf{x}(k),\epsilon_{\mathrm{c}}(k),\epsilon_{\mathrm{d}}(k))\\ \text{s.t. }&x(k+l+1)=Ax(k+l)+B_{1}u(k+l)+\\ &\hskip 28.45274pt+B_{2}\delta(k+l)+B_{3}z(k+l)+B_{5},\\ &E_{2}\delta(k+l)+E_{3}z(k+l)\leq E_{1}u(k+l)+\\ &\hskip 85.35826pt+E_{4}x(k+l)+E_{5},\\ &\text{for }l=0,\dots,N_{\mathrm{p}}-1\end{split}$ (2)
@@
the prediction horizon is denoted by $N_{\mathrm{p}}$, the variable $l$ indexes the time along the prediction horizon, and the final cost is represented by $V_{\mathrm{f}}(\cdot)$. The predicted state trajectory over the prediction horizon is denoted by $\mathbf{x}(k)=[x^{\mathrm{T}}(k),\ldots,x^{\mathrm{T}}(k+N_{\mathrm{p}})]^{\mathrm{T}}$. Let the predicted input be explicitly divided into continuous and discrete components $u(k)=[u_{\mathrm{c}}^{\mathrm{T}}(k),\ u_{\mathrm{d}}^{\mathrm{T}}(k)]^{\mathrm{T}}$. The stacked continuous input and auxiliary variables over the prediction horizon are represented in vector $\epsilon_{\mathrm{c}}(k)=[u_{\mathrm{c}}^{\mathrm{T}}(k),z^{\mathrm{T}}(k),\ldots,u_{\mathrm{c}}^{\mathrm{T}}(k+N_{\mathrm{p}}-1),\ z^{\mathrm{T}}(k+N_{\mathrm{p}}-1)]^{\mathrm{T}}$. Similarly, the stacked discrete inputs and auxiliary variables over the prediction horizon are expressed by $\epsilon_{\mathrm{d}}(k)=[u_{\mathrm{d}}^{\mathrm{T}}(k),\delta^{\mathrm{T}}(k),\ldots,u_{\mathrm{d}}^{\mathrm{T}}(k+N_{\mathrm{p}}-1),\ \delta^{\mathrm{T}}(k+N_{\mathrm{p}}-1)]^{\mathrm{T}}$. For the sake of simplicity, the constraints are assumed to be polyhedral as in (1).
At each time step, the optimization (2) is solved and the first entry of the decision variables is applied to the system. Let $(\mathbf{x}^{*}(k),\ \epsilon_{\mathrm{c}}^{*}(k),\ \epsilon_{\mathrm{d}}^{*}(k))$ be the optimal solution of (2). The MPC control law then is defined by
@@
If the cost function (3) is quadratic (linear), then the optimization problem (2) can be conveniently recast as a mixed integer quadratic (linear) program [4]. The corresponding quadratic optimization problem in a compact form can be stated as follows:
$\begin{split}\min_{\epsilon(k)}\ &\epsilon^{\mathrm{T}}(k)H_{1}\epsilon(k)+\epsilon^{\mathrm{T}}(k)H_{2}x(k)+c^{\mathrm{T}}\epsilon(k)\\ \text{subj. to }&G\epsilon(k)\leq w+Sx(k)\end{split}$ (5)
where $\epsilon(k)=[\epsilon_{\mathrm{c}}^{\mathrm{T}}(k),\ \epsilon_{\mathrm{d}}^{\mathrm{T}}(k)]^{\mathrm{T}}$, and $H_{1},\ H_{2},\ c,\ G,\ w,\ S$ are matrices of appropriate sizes. If the cost function is linear, then the matrices $H_{1}$ and $H_{2}$ are zero.

### v2

# Article Category

@@ Abstract
This work proposes an approach that integrates reinforcement learning and model predictive control (MPC) to solve finite-horizon optimal control problems in mixed-logical dynamical systems efficiently. Optimization-based control of such systems with discrete and continuous decision variables entails the online solution of mixed-integer linear programs, which suffer from the curse of dimensionality. Our approach aims to mitigate this issue by decoupling the decision on the discrete variables from the decision on the continuous variables. In the proposed approach, reinforcement learning determines the discrete decision variables and simplifies the online optimization problem of the MPC controller from a mixed-integer linear program to a linear program, significantly reducing the computational time. A fundamental contribution of this work is the definition of the decoupled Q-function, which plays a crucial role in making the learning problem tractable in a combinatorial action space. We motivate the use of recurrent neural networks to approximate the decoupled Q-function and show how they can be employed in a reinforcement learning setting. Simulation experiments on a microgrid system using real-world data demonstrate that the proposed method substantially reduces the online computation time of MPC while maintaining high feasibility and low suboptimality.
{IEEEkeywords}
learning for control, reinforcement learning, hybrid systems, predictive control for nonlinear systems, energy and power systems
@@ 1.1 Motivation
Complex infrastructure systems, such as energy, transportation, and water networks, are pervasive in our modern world. Analysis and control design of such systems is very challenging due to their size and intricate behavior. Modeling of such systems often require the combination of discrete and continuous decision variables to adequately capture the system dynamics and operational constraints. In this context, several modeling approaches [1, 2, 3] have considered hybrid systems to represent these critical infrastructure networks. For instance, the solution of the optimal dispatch problem in energy networks consists of scheduling both the power flow directions and intensities among each of the agents of the network. Some operational constraints require the addition of binary variables for modeling, see [1]. As a result, discrete and continuous decision variables have to be planned over a future horizon to reduce the operation cost considering market conditions.
@@
For the control of hybrid systems, model predictive control (MPC) arises as a promising technique due to its ability to handle constrained complex systems with discrete and continuous decision variables [4]. Furthermore, MPC has solid theoretical foundations with regard to stability, performance, and safety [5]. However, these advantages often come at the price of intense online computational requirements, limiting the use of MPC due to hardware constraints or execution time limitations. This bottleneck is especially more pronounced in MPC for hybrid systems, where the optimizer has also to consider a sequence of discrete decision variables over the given prediction horizon. In this case, the optimization problem then becomes mixed-integer program, which is NP-hard [6].
@@
The most widely used technique to solve mixed-integer program in modern solvers is branch-and-bound [7]. In the worst case, the solver has to find solutions for the relaxed problems for each possible combination of the discrete decision variables. This is not scalable with respect to the number of discrete variables because of the combinatorial nature of the problem. Branch-and-bound algorithms mitigate this issue by efficiently pruning branches of the search tree by estimating lower and upper bounds of the objective function and by proposing cutting planes, which reduce the feasible set of the mixed-integer programs by introducing linear inequalities as additional constraints [8]. Moreover, many expert-designed heuristics are employed for improving the search, such as node selection and branch variable selection techniques to reduce computation time. One alternative to branch-and-bound for control of hybrid systems is to pre-compute the MPC control law offline and simply evaluate this function online. For hybrid systems, the offline computation of the explicit MPC control law via multi-parametric programming and dynamic programming was explored in [6]. However, this approach, also referred to as explicit MPC, can be only successfully applied to low-dimensional linear systems. Despite improvements in branch-and-bound and explicit MPC, solving mixed-integer programs with a significant number of integer decision variables remains fundamentally difficult.
@@
Recently, supervised learning has been explored in several MPC approaches for hybrid systems to reduce the online computational time of the resulting mixed-integer programs [9, 10, 11, 12]. In essence, the aforementioned supervised learning methods have the same structure and learning setting and the main difference lies in the choice of the classifier. These works employ supervised learning to approximate the mapping from the system state to the discrete optimal solution. By applying this classifier to predict the discrete optimization variables, the mixed-integer program is then simplified to an optimization problem consisting only of real-valued variables, significantly reducing the online computational burden of the MPC controller. To build the training dataset, these methods rely on branch-and-bound to solve the control problem – a mixed-integer program – to optimality several times. As a result, the main computational issue is sidelined to the offline phase of the algorithm, where typically more computing resources are available. Moreover, in supervised learning, the goal is to reduce the classification error, e.g., the distance between the predicted discrete sequence and the optimal discrete sequence. Even though the classification error is typically a suitable proxy for control performance, this might not necessarily be true depending on the mapping from the input sequence to the control objective function. For a more complete overview of the literature on the intersection of learning and control for hybrid systems, the reader is referred to Section 2.
@@ 1.2 Contributions
We propose an integrated reinforcement learning (RL) and MPC method that solves mixed-integer linear programs with low computational footprint, low optimality gap, and high feasibility rate. We build on the existing idea of decoupling the decision on the discrete and continuous variables with learning and MPC [9, 10, 11, 12]. However, we explore a novel paradigm by employing reinforcement learning in place of supervised learning to directly optimize for control performance – instead of minimizing the classification error – and to avoid the use of branch-and-bound in both the offline and online phases of the algorithm.
@@
  • We propose a novel integrated reinforcement learning and MPC framework for control of mixed-logical dynamical systems. During online operation, the reinforcement learning agent simplifies a mixed-integer linear program into a linear program by fixing the discrete decision variables.
  • The Q-function is partitioned across the prediction horizon, and the definition of decoupled Q-functions is conceived to make the learning problem tractable. The decoupled Q-function is approximated by a recurrent neural network, and its role in the reinforcement learning algorithm is described.
  • Simulation experiments in a microgrid system show the efficacy of the proposed approach in reducing the computational load of the MPC controller. Moreover, the comparison between the proposed approach and a method based on supervised learning reveals a trade-off in the case study: while the former outperforms in terms of feasibility, the latter outperforms in terms of optimality.
@@ 1.3 Outline
This paper is organized as follows. In Section 2, we give an overview of the literature on the intersection of learning and control for hybrid systems. Section 3 formalizes the control problem. Our novel method that integrates reinforcement learning and MPC for control of mixed-logical dynamical systems is described in Section 4. Section 5 presents the simulation setup, the results, and the discussion. This paper ends in Section 6 with conclusions and suggestions for future work.
@@ 2 RELATED WORK
In [16] the state-action value function (Q-function) of time-invariant mixed-logical dynamics systems is approximated. The core of their approach is to parametrize the Q-function with Benders cuts and estimate the Q-function from a lower bound. However, the scope of application of this approach is limited because the paper does not consider discrete states and inputs. In [17], the computational burden of MPC is reduced by learning a state-dependent horizon and a state-dependent recomputation policy, i.e., the policy that decides whether the MPC optimal control problem should be recomputed at a given time step. An MPC controller is used as a function approximator of the state-action value in [18], where the parameters of the controller are tuned by policy gradient methods. An extension of the same approach to mixed-integer problems was made in [19]; however, the goal is to target performance rather than reduction in the computational cost of the optimization problem.
There is also a body of work that explores the use of reinforcement learning to jointly learn discrete and continuous policies [20, 21, 22]. In the artificial intelligence community, this problem is explored under the framework of Markov decision processes with parameterized actions. The problem is very similar to that of control of hybrid systems due to the nature of the action space, which has discrete and continuous elements. Even though these works can offer useful insight into the parametrization and training of policies for hybrid systems, they lack the optimality and constraint satisfaction that MPC can provide.
@@
In the literature regarding energy systems, several approaches exist for the integration of learning and optimization-based control. In [25], a parameterized MPC controller is tuned by an actor-critic reinforcement learning method to improve the performance of a home energy management system. A similar approach is developed in [26] for a residential microgrid system. Therein, in addition to the parameterized MPC approach, the energy management problem is framed as a cooperative coalition game, and the Shapley value is used to distribute the costs between the consumers. The authors of [27] propose the MIP-DQL algorithm in which the Q-function maximization problem is represented as a mixed-integer program, in which the system constraints are embedded. In this fashion, the reinforcement learning algorithm can guarantee operational constraints via online optimization. Another use of learning is to estimate prediction models for the MPC controller. In [28], grey-box and neural network models are compared for multistep-ahead prediction for control of heating, ventilation, and air-conditioning systems. The paper [29] proposes the use of learning to assist column generation in the solution of the mixed-integer programs with the aim of reducing the computational cost of MPC-based approaches in energy management for microgrids.
The use of mixed-logical dynamical systems for modeling and control of microgrids has been considered in [1], [30], [10], and [9]. In these works, an MLD system is employed to cast a mixed-integer linear problem for microgrid operation optimization. A similar model is used in the case study of our work to solve the optimal dispatch problem.
@@ 3 CONTROL PROBLEM
We consider MPC for its capacity to handle multivariable constrained hybrid systems. In this context, mixed-logical dynamical (MLD) systems [31] are typically used to formulate open-loop finite-horizon optimal control problems [4]. Moreover, the equivalence of MLD systems and other hybrid system modeling frameworks was established in [32], showing their broad applicability. When MLD systems are used to formulate the MPC problem and the cost function and the constraints are linear, the resulting optimization problem is a mixed-integer linear program (MILP). In this section, a general description of the MPC optimization problem for an MLD system is first addressed to set the stage for the formulation of the control problem as an MILP.
@@
where $x\in\mathbb{R}^{n_{\mathrm{c}}}\times\{0,1\}^{n_{\mathrm{d}}}$ is a vector containing the continuous and discrete system states, $u\in\mathbb{R}^{m_{\mathrm{c}}}\times\{0,1\}^{m_{\mathrm{d}}}$ are the continuous and discrete inputs, $\delta\in\{0,1\}^{r_{\mathrm{d}}}$ is a vector with the auxiliary discrete variables, $z\in\mathbb{R}^{r_{\mathrm{c}}}$ is a vector with the continuous auxiliary variables arising from the MLD modeling, and $A,\ \{B_{i}\}_{i=1,2,3,5},\ \{E_{i}\}_{i=1,2,3,4,5}$ are matrices of appropriate dimensions. Note that the linear constraints may represent logical constraints, a byproduct of MLD modeling, or system operating constraints. For a demonstration of the application of the discrete variables and auxiliary variables to represent logical constraints, the reader is referred to Appendix 7, where the model for the case study is characterized.
@@
$\begin{split}\min_{\mathbf{x}(k),\epsilon_{\mathrm{c}}(k),\epsilon_{\mathrm{d}}(k)}\ &J(\mathbf{x}(k),\epsilon_{\mathrm{c}}(k),\epsilon_{\mathrm{d}}(k))\\ \text{s.t. }&x(k+l+1)=Ax(k+l)+B_{1}u(k+l)+\\ &\hskip 28.45274pt+B_{2}\delta(k+l)+B_{3}z(k+l)+B_{5},\\ &E_{2}\delta(k+l)+E_{3}z(k+l)\leq E_{1}u(k+l)+\\ &\hskip 85.35826pt+E_{4}x(k+l)+E_{5},\\ &\text{for }l=0,\dots,N_{\mathrm{p}}-1,\end{split}$ (2)
@@
where the prediction horizon is denoted by $N_{\mathrm{p}}$, the variable $l$ indexes the time along the prediction horizon, and the final cost is represented by $V_{\mathrm{f}}(\cdot)$. The predicted state trajectory over the prediction horizon is denoted by $\mathbf{x}(k)=[x^{\mathrm{T}}(k),\ldots,x^{\mathrm{T}}(k+N_{\mathrm{p}})]^{\mathrm{T}}$. Let the predicted input be explicitly divided into continuous and discrete components $u(k)=[u_{\mathrm{c}}^{\mathrm{T}}(k),\ u_{\mathrm{d}}^{\mathrm{T}}(k)]^{\mathrm{T}}$. The stacked continuous input and auxiliary variables over the prediction horizon are represented in vector $\epsilon_{\mathrm{c}}(k)=[u_{\mathrm{c}}^{\mathrm{T}}(k),z^{\mathrm{T}}(k),\ldots,u_{\mathrm{c}}^{\mathrm{T}}(k+N_{\mathrm{p}}-1),\ z^{\mathrm{T}}(k+N_{\mathrm{p}}-1)]^{\mathrm{T}}$. Similarly, the stacked discrete inputs and auxiliary variables over the prediction horizon are expressed by $\epsilon_{\mathrm{d}}(k)=[u_{\mathrm{d}}^{\mathrm{T}}(k),\delta^{\mathrm{T}}(k),\ldots,u_{\mathrm{d}}^{\mathrm{T}}(k+N_{\mathrm{p}}-1),\ \delta^{\mathrm{T}}(k+N_{\mathrm{p}}-1)]^{\mathrm{T}}$. For the sake of simplicity, the constraints are assumed to be polyhedral, as in (1).
At each time step, the optimization (2) is solved, and the first entry of the decision variables is applied to the system. Let $(\mathbf{x}^{*}(k),\ \epsilon_{\mathrm{c}}^{*}(k),\ \epsilon_{\mathrm{d}}^{*}(k))$ be the optimal solution of (2). The MPC control law then is defined by
@@
If the cost function (3) is linear, e.g., state and terminal costs based on the 1-norm or the $\infty$-norm, then the optimization problem (2) can be conveniently recast as a mixed integer linear program [4]. The corresponding linear optimization problem in a compact form can be stated as follows:
$\begin{split}\min_{\epsilon(k)}\ &c^{\mathrm{T}}\epsilon(k)\\ \text{s.t. }&G\epsilon(k)\leq w+Sx(k)\end{split}$ (5)
where $\epsilon(k)=[\epsilon_{\mathrm{c}}^{\mathrm{T}}(k),\ \epsilon_{\mathrm{d}}^{\mathrm{T}}(k)]^{\mathrm{T}}$, and $c,\ G,\ w,\ S$ are matrices of appropriate sizes.
Note that the optimization problem (5) is parameterized by the state $x(k)$. Besides, the system state has an effect on the objective through the auxiliary variable $z(k)$, which is part of the decision variable $\epsilon(k)$. Similarly, the resulting optimization problem can also possibly depend on an exogenous variable $\gamma(k)$, i.e., a measured external variable outside the influence of the controller. The effect of $\gamma(k)$ on the optimization problem is to further parameterize the mixed-integer program. In principle, the exogenous variable can potentially change any of the matrices $c,\ G,\ w,$ and $S$ that define (5). In our problem formulation, we restrict the parametrization to the cost vectors $c$ and the constraint matrix $w$.
The problem (5) can then be extended to reflect the addition of the exogenous variable:
$\begin{split}\min_{\epsilon(k)}\ &c^{\mathrm{T}}(\gamma(k))\epsilon(k)\\ \text{s.t. }&G\epsilon(k)\leq w(\gamma(k))+Sx(k)\end{split}$ (6)
For ease of notation, we leave out the explicit dependence of such parameters on the exogenous signal hereafter. Since (6) is parameterized both by the system state $x(k)$ and the exogenous signal $\gamma(k)$, we define the augmented state
$\chi(k)=[x^{\mathrm{T}}(k),\ \gamma^{\mathrm{T}}(k)]^{\mathrm{T}}$
to fully characterize the parameters that define the parametric mixed-integer linear program.
The exogenous variables are necessary for the modeling of some applications, as shown in the case study in Section 5, where such variables are comprised of electricity prices, renewable energy generation, and energy demand forecasts.
Remark.
The assumption on the linearity of the objective function does not play any fundamental role in the development of the proposed approach in the next section. The choice for a linear objective is to simplify the notation and the presentation of this work. In principle, even a nonconvex objective cost could be chosen. Such a choice would, however, result in more complex optimization problems to be solved offline for learning and solved online for inference. Similarly, even though (6) can be formulated in our proposed approach with a more general parametrization in the exogenous variable $\gamma(k)$, we restrict the parametrization for the sake of simplicity and alignment with the case study.